OpenStudy (anonymous):

assume x=x(t) and y=y(t). Find dx/dt if x^2(y-6) = 12y + 3 and dy/dt = 2 when x = 5 and y = 12

OpenStudy (anonymous):

I'm pretty lost in general about this problem but I think it's supposed to use differential equations!

OpenStudy (dls):

Simply differentiate it to get \[\Large 2x \frac{dx}{dt}(y-6) + x^2 (\frac{dy}{dt})=12 \frac{dy}{dt}\] solve for dx/dt by rearranging terms x=5,y=12 dy/dt=2

OpenStudy (anonymous):

So when you say solve for dx/dt by rearranging, would it be: \[(12-6)+5 ^{2}(2)-12(2)=2(5)\frac{ dx }{ dt } ? \] Or is that incorrect? I feel 100% sure that I made an error somewhere but I'm not sure how to separate it honestly.

OpenStudy (dls):

\[\Large \frac{dx}{dt}=\frac{12 \frac{dy}{dt}-x^2 \frac{dy}{dt}}{2x}\]

OpenStudy (dls):

now put all the values

OpenStudy (anonymous):

\[\frac{ dx }{ dt }= \frac{ 24 - 50 }{ 10 } = \frac{ -26 }{ 10 }= \frac{ -13 }{ 5 }\] Is that correct?

OpenStudy (anonymous):

or should it be 24+50

OpenStudy (dls):

-13/5 is absolutely correct mate :) well done

OpenStudy (anonymous):

Here's the confusing part though. The only viable answer choices on my test (I've already taken it btw) are: \[-\frac{ 20 }{ 13 } | -\frac{ 13 }{ 30 } | \frac{ 13 }{ 20 } | \frac{ 20 }{ 13 }\] I circled the third option 13/20 and got the question wrong and -13/5 doesn't match any solution unfortunately 8(

OpenStudy (anonymous):

Anyone have an idea where I went wrong?

OpenStudy (anonymous):

or what went wrong?

ganeshie8 (ganeshie8):

x^2(y-6) = 12y + 3 2x(y-6) + x^y' = 12y' y' = (2x(y-6))/(12-y^2) y' at (5, 12) = -60/13

ganeshie8 (ganeshie8):

dy/dx = dy/dt / dx/dt -60/13 = 2/A solve A

OpenStudy (anonymous):

-13/30! I understand it now! Thanks !

ganeshie8 (ganeshie8):

np :) im still wondering why the other method was not giving the same answer :/